OpenStudy (anonymous):

given the function g described by g(x)=x+5 find the function value g(c+7)

OpenStudy (anonymous):

wherever you see an x, replace that with c+7: g(c+7) = (c+7) + 5 = c + 12

OpenStudy (anonymous):

Yep, just plug in!